Solve the following non-homogeneous linear ODE of first order

dy/dx + 3y/3x = 6x^2

 



1
Expert's answer
2021-04-15T16:57:36-0400

Given,

"\dfrac{dy}{dx}+\dfrac{3y}{3x}=6x^2"


This is linear differential equation so


Here "P=\dfrac{1}{x}, Q=6x^2"


Integrating factor "I.F.=e^{\int Pdx}=e^{\dfrac{1}{x}dx}=e^{logx}=x"


Therefore the solution is-


"y\times I.F.=\int Q\times I.F. dx"


"y\times x=\int 6x^3dx"


"y\times x=\dfrac{6x^4}{4}+C"


"yx=\dfrac{3}{2}x^4+C"


Hence, "y=\dfrac{3}{2}x^3+\dfrac{C}{x}"